Large pizzas cost $18.00 and small pizzas cost $11.00. The restaurant sold 34 pizzas with a total value of $479. How many of each did they sell?
Answer by Cromlix(4381)

Large pizzas = x
Small pizzas = y
x + y = 34...1
18x + 11y = 479...2
Multiply 1 by 11
11x + 11y = 374...3
18x + 11y = 479...2
Subtract (3) from (2)
7x = 105
x = 15
Substitute x = 15 into (1)
x + y = 34
15 + y = 34
y = 34 - 15
y = 19
So, 15 large and 19 small pizzas were sold.
Hope this helps.
:-)